OpenStudy (anonymous):

Write in equation in slope intercept form for the line that passes through (-1,-2) and (3,4)

OpenStudy (shamil98):

Do you know how to find slope?

OpenStudy (shamil98):

\[\large m =\frac{ \Delta y }{ \Delta x } = \frac{ y_2-y_1 }{ x_2-x1 }\]

OpenStudy (anonymous):

Subtract y that goes on the top then you subtract x

OpenStudy (shamil98):

Find your slope first.

OpenStudy (anonymous):

Is the slope 6/ 4

OpenStudy (shamil98):

4 - - 2 / 3 - - 1 = 6/4 yes. but you can reduce it to 3/2

OpenStudy (anonymous):

Okay

OpenStudy (shamil98):

From, there input anyone of your group of coordinates into y = mx + b and solve for b.

OpenStudy (shamil98):

Let's use (3,4) 4 = 3/2(3) + b 4 = 9/2 + b 4 - 9 /2 = b -1/2 = b y = (3/2)x - 1/2 is your slope-intercept form

OpenStudy (anonymous):

Thx